Rewrite the equation so that x is a function of y and then use the result to find x when y =-2
4(5-y)=14x+3

Found 2 solutions by stanbon, checkley77:
Answer by stanbon(26284)

Rewrite the equation so that x is a function of y and then use the result to find x when y =-2
4(5-y)=14x+3
20-4y = 14x + 3
-4y = 14x - 17
y = (-14/4)x + (17/4)
-----------------------------
If y = -2 solve for "x":
4(5-y) = 14x+3
4(7) = 14x+3
25 = 14x
x = 25/14
---------------
Cheers,
Stan H.

Answer by checkley77(7068)

4(5-y)=14x+3
20-4y=14x+3
-4y=14x+3-20
-4y=14x-17
y=14x/-4-17/-4
y=-7x/2+17/4
----------------------
-2=-7x/2+17/4
-7x/2=-2-17/4
-7x/2=(4*-2-17)/4
-7x=2(-8-17)/4
-7x*4=2*-25
-28x=-50
x=-50/-28
x= 3/25/142 ans.
Proof:
-2=-7/2*25/14+17/4
-2=-175/28+17/4
-2=(-175+17*7)/28
-2=(-175+119)/28
-2=-56/28
-2=-2